OpenStudy (anonymous):

If f(g(x) )=6x+1 and f(x)=2x+3, what is g(x)? Explain the answer, please. Thanks! :))

OpenStudy (lgbasallote):

first let me ask you...say f(x = 2x + 3...how would you solve for f(6)?

OpenStudy (lgbasallote):

note..im asking HOW..not WHAT

OpenStudy (anonymous):

by substitution. that is f(6) = 2(6) + 3 = 15

OpenStudy (lgbasallote):

good! f(g(x)) is the same...you substitute g(x) to every x term in f(x) 2x + 3 only x term is in 2x + 3 f(g(x)) = 2(g(x)) + 3 got it?

OpenStudy (anonymous):

2 g(x) + 3 = 6x + 1 -> g(x) = ....

OpenStudy (lgbasallote):

^that's the continuation

OpenStudy (anonymous):

@lgbasallote clarify the relation :)

OpenStudy (lgbasallote):

will do next time :D

OpenStudy (anonymous):

thanks @lgbasallote and @Chlorophyll . got the answer, g(x)= 3x -1. :))

OpenStudy (anonymous):

:)